OpenStudy (anonymous):

Write a linear equation to model the situation. You borrow $70 from your brother. To repay the loan, you pay him $7 per week. Use w to denote the number of weeks and y as the balance on the loan.

OpenStudy (mathstudent55):

You start by owing $70. Since y is the balance of the loan, you start with y = 70. Each week, you pay $7. After the first week, w = 1, you have paid 7 * 1. Now the balance is y = 70 - 7 After the second week, w = 2, you have paid 7 * 2. Now the balance is y = 70 - 14 After week w, w = w, you have paid 7 * w = 7w. Now the balance is ? Can you finish it?

OpenStudy (anonymous):

Yes. Thank you for your help.

OpenStudy (mathstudent55):

You're welcome.